OpenStudy (anonymous):

Help, please? I'm confused. xD The graph of a system of equations will intersect at more than 1 point.

OpenStudy (anonymous):

Answer options are: Sometimes Never Always

OpenStudy (anonymous):

I'm pretty sure it might be Never..

jimthompson5910 (jim_thompson5910):

If you're talking about linear equations, then they will only intersect more than once only if you have a dependent system. So this happens some of the time but not all the time. Ex: 2x+3y = 6 4x+6y = 12 is a dependent system that has an infinite number of intersections

OpenStudy (anonymous):

the answer is sometimes

OpenStudy (anonymous):

Now I'm really confused. ._.

jimthompson5910 (jim_thompson5910):

The lines 2x+3y = 6 and 4x+6y = 12 lie on top of one another (since they're the same line). So they intersect at an infinite number of points.

OpenStudy (anonymous):

OH. I see what you mean. n.n Thanks.

jimthompson5910 (jim_thompson5910):

That's just one example of a system that has more than one intersection

jimthompson5910 (jim_thompson5910):

you're welcome